Hybridization of a homozygous dominant and homozygous recessive plant was conducted. What is the probability that the F2 will have the phenotype big oval shaped yellow leaves, white flowers, and pink fruits?
1
Expert's answer
2020-11-19T09:48:24-0500

Suppose that all alleles in genes are in a relationship of complete dominance, then we can assume:

A-Oblong leaves

a-Oval leaves

B-Green leaves

b- Yellow leaves

C- Purple flowers

c- White flowers

D- Red fruits

D-Pink fruits

F2- aabbccdd (phenotype big oval shaped yellow leaves, white flowers, and pink fruits)?

P: AABBCCDD x aabbccdd

F1:    AaBbCcDd

To analyze phenotypes, we will use the most technically simple mathematical method.

Mathematical (algebraic) method - the probability of occurrence of any combination of all genotypes in a polyhybrid crossing is equal to the product of the probabilities of genotypes for each of the analyzed genes.

At interbreeding: P AaBbCcDd x AaBbCcDd

Crossing for gene A: Аа × Аа;

F2 for gene A: (¾ А_ + ¼аа)

 

Crossing for gene В: Вb × Вb

F2 for gene В: (¾ В_ + ¼bb)


Crossing for gene C: Cc × Cc

F2 for gene C: (¾ C_ + ¼cc)

 

Crossing for gene D: Dd × Dd

F2 for gene D: (¾ D_ + ¼dd)

 

For definition in F2 aabbccdd is necessary:

aabbccdd = (¼аа) * (¼bb) * (¼cc) * (¼dd) = 1/256 = 0,004%